OpenStudy (anonymous):

Simplify (3+3i)^5 using exponential form and express in a+bi form.

jhonyy9 (jhonyy9):

us formula (a+b)^5 =(a+b)^2 *(a+b)^3 but attention !!! on i^2 = -1 !!!!!